OpenStudy (anonymous):

Solve for x: (3/2)+(1/4)x = (1/6)-(5/3)x

OpenStudy (anonymous):

(5/3)x+(1/4)x = (1/6) -(3/2) ===> (23/12)x = -4/3 ==> x= -23/16

OpenStudy (dumbcow):

Move the "x" terms to one side and move the constants all to one side --> (1/4)x +(5/3)x = (1/6) -(3/2)

OpenStudy (anonymous):

i got -16/23.

OpenStudy (dumbcow):

you are correct, its -16/23